) /_PQR and /_ABC are complementary angles and /_PQR is nine times as large as /_ ABC. Determine the measure of each angle.

Answer by checkley77(12844)

The sumof two complementary angles=90 degrees.
PQR=9ABC
PQR+ABC=90
9ABC+ABC=90
10ABC=90
ABC=90/10
ABC=9 DEGREES.
PQR=9*9
PQR=81 DEGREES.
PROOF:
9+81=90